OpenStudy (anonymous):

Please Help me! What is the slope of the line that passes through the pair of points? (1, 7), (10, 1)

OpenStudy (anonymous):

ok so for this question your going to have to do linear equations and parabolas... do you know how to do that?

OpenStudy (anonymous):

np i do not..

OpenStudy (anonymous):

ok well get a piece of graph paper to do it

OpenStudy (anonymous):

now you have to make the parabola

OpenStudy (anonymous):

to make the parabola you make an x and y chart so put under x 1 and y 7. graph the x and y on the graph

OpenStudy (anonymous):

let me get a graph paper quick...

OpenStudy (anonymous):

slope m = (Y2 - Y1) / (X2 - X1) = (1 - 7 )/ (10 - 1) =

OpenStudy (anonymous):

-6 / 9 => -2 / 3

OpenStudy (anonymous):

now graph the other points x 10 and y 1

OpenStudy (anonymous):

make an equation, connect the lines, add the parabola by substituting for x and getting your y and mark it as the curve line. so youve got your two points which are... can you figure this out?